4. The following income distribution data are for a hypothetical developing country. (20 points) Ọuintile Lowest 20% Second quintile Third Fourth quintile Highest 20% Highest 10% Percent Share 2.4% 5.7% 10.7% 18.6% 62.6% 46.2% quintile a) Carefully graph the Lorenz curve, labeling the axes. (5 points) b) This countrys national income is about $300 billion. What is the dollar income of the bottom 20 percent? Bottom 40 percent? (5 points) c) Using the same income data specified in (ii), assume that this countrys population is 150 million. Suppose that each household earns the average income for its quintile. How many people in this country are poor if the poverty line is S400 per capita? How much will it take to lift the poor out of poverty? (10 points)
Answer #1

a)

Population

Perfect equality distribution

Actual Income

Cumulative distribution

0%

0%

0.00%

0.00%

20%

20%

2.40%

2.40%

40%

40%

5.70%

8.10%

60%

60%

10.70%

18.80%

80%

80%

18.60%

37.40%

100%

100%

62.60%

100.00%

w88BPW1mnSQGAAAAABJRU5ErkJggg==

b)

National Income=NI=$300 billion

Dollar income of 20% population =2.40% of NI=300*2.4%=$7.2 billion

Dollar income of 40% population =8.10% of NI=300*8.1%=$24.3 billion

c)

Population

Population, million

Actual Income f

$ income= f*300000 million

Income per capita

Lowest 20%

30

2.40%

7200

240

Second Quintile

30

5.70%

17100

570

Third Quintile

30

10.70%

32100

1070

Fourth Quintile

30

18.60%

55800

1860

Highest 20%

30

62.60%

187800

6260

We find that Lowest 20% population is below poverty level income of $400 per capita.

No. of persons in this category are 30 million.

Amount needed to uplift this population=(400-240)*30=$4800 million or say $4.8 billion